A bike originally costs $185.99, however, there is a 30% off sale. Then there is a scratch on the bike, so
they are willing to take an additional 10% off the sale price. With 13% hst tax, what is the final price of
the bike?

Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39618)

this was previously answered.

Answer by ikleyn(52788)

.

We have this chain of multiplicative coefficients to apply to the original price 

        1-0.3 = 0.7;  1-0.1 = 0.9  and  1+0.13 = 1.13.


Therefore, the ANSWER to the problem's question is


        0.7*0.9*1.13*185.99 = 132.4063 = 132.41  cents  (rounded to the nearest cent).

Solved.

-------------

See the lesson
    - Percentage problems on chains of discounts
in this site.


/\/\/\/\/\/\/\/

For clarity:  I placed my solution here because the solution by @josgarithmetic is  INCORRECT.